# SolNet: Open-source deep learning models for photovoltaic power forecasting across the globe

### v2

@@ 4.5 The effect of source data mis-specification
Two key findings become clear from these figures. On the one hand, we can see that mis-specification of the target location comes with increasingly erroneous forecasts. On the other hand, the use of external covariates plays a key role in the impact of these mis-specifications. When weather covariates are included in the model (Figure 5(b)), the results worsen drastically with increased distance from the target location. Moving 100km to the south-east increases the error by 15% (from 0.080 to 0.093) over the correct location, 35% at 200km (from 0.080 to 0.108) and at 400km the error is already at a similar level to a naive model. When no weather variables are included (Figure 5(a)), this effect is far less pronounced. The performance definitely lags behind on a model with weather covariates, but the model is much more robust towards mis-specifications.
The median target RMSE (the blue line) also makes clear how important correct specification of the target location is when weather data is included. When the location is mis-specified by c. 200km we can see that the transfer model has already lost all its relative value after two months. When the mis-specification is as small as 100km, the value is lost after 5 months. However, when no weather data is included, a mis-specification of the source domain has almost no effect on target outperformance, even after more than 6 months.
@@
Figure 5: Transfer learning results under increasing distance from the target for a subset of ten NL locations. The performance is compared to the median baseline of these ten locations, as well as the median ’correct’ target model. I.e. a target model that is made with no mis-specification.
@@ 5 Discussion
There are several important considerations to be made to assess the value of transfer learning. Weather data has a strong positive effect on the predictive performance of models, and the translation from actuals in the source model as a proxy for forecasts, to actual forecasts holds up. If no data is available from a sunnier time period, the transfer model will have considerably more added value that will last longer, although here as well, weather data in the transfer model mitigates this effect. This indicates that most of the difficulty is in correctly learning to forecast the seasonality inherent in PV energy production and, more importantly, the peak production. Finally, it is important to try to mimic the situation of the target data (location, azimuth, slope) to limit the distance between source and target domain. However, the transfer models do appear to be robust to small mis-specifications, but larger mis-specifications can have a detrimental effect, especially when detailed external covariates, tied with the location, are used for predictive purposes.
@@ 6 Conclusion
This paper provides a detailed assessment of the impact of utilising a large corpus of open-source (synthetic) data in the creation of data-driven forecast models for solar power output. The improvements demonstrated hold regardless of the location of the PV panels, and are rather robust to mis-specifications in system parameters. As evidence, the paper looked at over 300 sites in three different countries to substantiate the findings while running extensive tests lasting one year on all sites.
This methodology extends existing recent work on transfer learning by minimising the need for data from ’similar sites’, and shows that there is a large boost to the initial predictive performance. As expected, this differential subsides over time, and is almost zero after a one year period, but the impact of using transfer models is accumulative over time and dependent on several other factors like seasonality, giving users a powerful tool and a set of guidelines to improve their forecasting performance at the early onset of (or even before) installing a new PV setup.
The code, along with a functioning example for using SolNet for solar forecasting anywhere on land around the globe, has been open-sourced [^3] https://anonymous.4open.science/r/SolNet-CD07/README.md. We expect this work to be the first of many that have the potential to lead to general-purpose deep learning models for energy forecasting.

### v3

@@ 4.4 The effect of source data misspecification
Two key findings become clear from these figures. On the one hand, we can see that misspecification of the target location comes with increasingly erroneous forecasts. On the other hand, the use of external covariates plays a key role in the impact of these misspecifications. When weather covariates are included in the model (Figure 5(b)), the results worsen drastically with increased distance from the target location. Moving 100km to the south-east increases the error by 15% (from 0.080 to 0.093) over the correct location, 35% at 200km (from 0.080 to 0.108) and at 400km the error is already at a similar level to a naive model. When no weather variables are included (Figure 5(a)), this effect is far less pronounced. The performance definitely lags behind on a model with weather covariates, but the model is much more robust towards misspecifications. The effect on the MAE (figure 8) is similar to that of the RMSE. Weather data adds in predictive performance, but has a detrimental effect when the target domain is misspecified.
The median target RMSE (the blue line) also makes clear how important correct specification of the target location is when weather data is included. When the location is misspecified by c. 200km we can see that the transfer model has already lost all its relative value after two months. When the misspecification is as small as 100km, the value is lost after 5 months. However, when no weather data is included, a misspecification of the source domain has almost no effect on target outperformance, even after more than 6 months.
@@
Figure 5: Transfer learning results under increasing distance from the target for a subset of ten NL locations. The performance is compared to the median baseline of these ten locations, as well as the median ’correct’ target model. I.e. a target model that is made with no misspecification.
@@ 5 Discussion
There are several important considerations to be made to assess the value of transfer learning. Weather data has a strong positive effect on the predictive performance of models, and the translation from actuals in the source model as a proxy for forecasts, to actual forecasts holds up. If no data is available from a sunnier time period, the transfer model will have considerably more added value that will last longer, although here as well, weather data in the transfer model mitigates this effect. This indicates that most of the difficulty is in correctly learning to forecast the seasonality inherent in PV energy production and, more importantly, the peak production. Finally, it is important to try to mimic the situation of the target data (location, azimuth, slope) to limit the distance between source and target domain. However, the transfer models do appear to be robust to small misspecifications, but larger misspecifications can have a detrimental effect, especially when detailed external covariates, tied with the location, are used for predictive purposes.
@@ 6 Conclusion
This paper provides a detailed assessment of the impact of utilising a large corpus of open-source (synthetic) data in the creation of data-driven forecast models for solar power output. The improvements demonstrated hold regardless of the location of the PV panels, and are rather robust to misspecifications in system parameters. As evidence, the paper looked at over 300 sites in three different countries to substantiate the findings while running extensive tests lasting one year on all sites.
This methodology extends existing recent work on transfer learning by minimising the need for data from ’similar sites’, and shows that there is a large boost to the initial predictive performance. As expected, this differential subsides over time, and is almost zero after a one year period, opposed to what we hypothesised (see figure 1). But the impact of using transfer models is accumulative over time and dependent on several other factors like seasonality, giving users a powerful tool and a set of guidelines to improve their forecasting performance at the early onset of (or even before) installing a new PV setup.
The code, along with a functioning example for using SolNet for solar forecasting anywhere on land around the globe, has been open-sourced [^4] https://github.com/Joris-Dprtr/SolNet. We expect this work to be the first of many that have the potential to lead to general-purpose deep learning models for energy forecasting.
Appendix A Supplementary figures
(a) NL scaled MAE
(b) NL scaled MBE
(c) AUS scaled MAE
(d) AUS scaled MBE
(e) BE scaled MAE
(f) BE scaled MBE
Figure 6: The mean absolute error (MAE) and mean biased error (MBE) from the base experiments. Compared to the RMSE, the MAE of the target and transfer models is notably worse relative to the baseline. In terms of bias, a target made model suffers from positive bias until several months of data are included in the training process.
(a) NL scaled MAE including weather variables
(b) NL scaled MBE including weather variables
Figure 7: The mean absolute error (MAE) and mean biased error (MBE) from the experiments including weather data. Once weather data is included in the models, the MAE follows the same trend as the RMSE, with the transfer and the target model (eventually) outperforming the benchmark. The bias is also gone faster in a target model once weather data is included.
(a) NL scaled MAE with increasing distance from the target
(b) NL scaled MBE with increasing distance from the target
(c) NL scaled MAE with increasing distance from the target including weather variables
(d) NL scaled MBE with increasing distance from the target including weather variables
Figure 8: The mean absolute error (MAE) and mean biased error (MBE) from the experiments with increasing misspecification of the target domain. The top figures do not have weather data included and the bottom figures do have weather data included. Note that the figures without weather data no longer align with 6 (which is clear from the negative bias that can now be seen in the target model in 8(b)) as different months are used in order to compare the results with the results including weather data (for which only 8 months are available, instead of 12).